# CU-JADE: A Method for Traversing Extinction Jumps along the Line-of-Sight

### v1

@@ A.2 The “noise” of CUSUM
[…]
where $\mu$ is the absolute difference between the mean of each segment and the overall mean $\bar{A}$, $\sigma$ is the dispersion of $A_{\lambda}$ of each segment, and $n$ is the equivalent number of steps. This expression does not have a simple elementary function form, but it can be split into two parts and solved using the error function ($\rm erf$), which is a related function of the Gaussian integral.
@@
The intricate statistical relationship between $S^{0}_{\rm diff}$ and $S_{\rm diff}$ allows for more effective utilization of physical quantities like $\Delta A_{\lambda}$, $\sigma$, and $n$, aligning closely with outcomes from manual verification. Moreover, validations and implementations in this work demonstrate that the derived “noise” in CUSUM offers an efficient approach to assess the $CL$.
Figure A2: The $D$-$S$ plot in the presence of a slope in $D$-$A$ space. Six false change points are noted by red crosses (I, II, III, IV, V, VI are the serial numbers of the detection in iterations). The curve is best fit by using Least Squares method (in green). In comparison, a predicted parabola curve (in blue) is drawn using our preset parameters. The colored curves are the $D$-$S$ of shuffled samples.
@@ A.3 The False Jump Features from Slope in D-A Map
We assume the extinction of diffuse gas still follows a Gaussian distribution on the $D$-$A$ map, yet it is superimposed on a linear function. The CUSUM statistic $S$ satisfies
@@
In the case, we set $S\left(0\right)=0$. The location of the symmetry axis, and/or the first false change point to be detected, is in $x_{\rm I}=\frac{\left(\bar{Y}-m\right)}{k}+\frac{1}{2n}$. Because $\frac{1}{2n}$ is smaller than the distance resolution in a given number density $n$, we ignore it and $x_{\rm I}\approx\frac{\left(\bar{Y}-m\right)}{k}$. CUSUM $S$ can thus be wrote as
@@
As illustrated in Figure A2, the $D$-$S$ plot draws a parabolic trace as expected. In a smaller $CL$ threshold ($C$=1 in Mode I), six false change points were detected in order, all of which are located near the midpoints of their respective iteration intervals. In comparison, we give the best fit by using Least Squares method (green curve in Figure A2) and a predicted parabola curve from the preset parameters (blue curve in Figure A2) by Equation (A10). Stricter criterion is needed to avoid such false detections. Next, we have to further analyze the properties of “noise” $S_{\rm diff}$.
Figure A3: Relations between $S_{\rm diff}$ and $A_{\lambda}$ dispersion $\sigma$, number density $n$, and slope $k$. (a) Set $n=0.1$ ${\rm pc}^{-1}$ and $k=0.003$ ${\rm mag}\cdot{\rm pc}^{-1}$, $S^{\rm avg}_{\rm diff}\propto\sigma$ when $\sigma$ become larger. (b) $S^{\rm avg}_{\rm diff}\propto\sqrt{n}$. (c) Set $\sigma$=1 mag, the relation between $S^{\rm avg}_{\rm diff}$ and $k$ shows a similar trend to panel (a). Panel (d) is a zoom-in from panel (c).
Similar to the situation where there are jump points, the $S_{\rm diff}$ with a slope of $k$ and dispersion $\sigma$ follows a random walk with variable steps (see Appendix A.1). Providing an analytical form for $S_{\rm diff}$ is equally complex. Generally, $S_{\rm diff}$ is proportional to $\sqrt{n}$, while $k$ and $\sigma$ also present a competitive relationship. For example, when $k$ is small, $\sigma$ becomes dominant. And the relation between $S_{\rm diff}$, $n$, $k$, and $\sigma$ can also be derived using mock data. By examining the typical values from the $D$-$S$ diagram along the LoS, we can describe the relationship between $S_{\rm diff}$ and one of the quantities.
@@
However, we find that $\sigma$ domains $S_{\rm diff}$ in small $k$ values. Figure A3 (d) shows a zoomed-in view of the small $k$ region, where $S_{\rm diff}$ climbs slowly within the range of small $k$. In a simple test on the mock data, assuming the following preset parameters: a typical value of jumps intervals $x_{\rm I}=750$ pc, a gradient $k=0.002$ ${\rm mag}\cdot{\rm pc}^{-1}$, a stars density $n=0.1~{}{\rm pc}^{-1}$, and $A_{\lambda}$ dispersion $\sigma=1~{}{\rm mag}$, we compute the prominence of the preset slope. The calculated prominence is $S_{\rm diff}^{0}\sim 50$ and the averaged $S_{\rm diff}\sim 20$. It can effectively avoid false detections caused by slopes when setting $C\geq 2$, and the percentile to 0.95. Of course, for larger $\Delta A_{\lambda}$, we could raise the confidence threshold to ensure a more reliable result. On the other hand, if $k\gtrsim 0.003$ ${\rm mag}\cdot{\rm pc}^{-1}$, we can find these change points near midpoint of each iteration, due to the symmetric/parabolic nature of $S_{\rm diff}$. This may help us determine if a extinction structure is from a real jump or a large slope $k$ (see Figure B1 in Appendix B).
@@ Appendix C Examples for HMSFRs associated with masers
Refer to caption Refer to caption
(a) G079.87+01.17 (b) G074.56+00.84
Refer to caption Refer to caption
(c) G012.68-00.18 (d) G209.00-19.38
Figure C1: Four examples for G079.87+01.17, G074.56+00.84, G012.68-00.18, and G209.00-19.38. Stars are represented by grey dots, while black dots denote binned points every 20 pc. All four examples corresponding to the masers are located at the last jump point.
In Section 4.2, we found that a total of 46 masers align well with our results. Here we provide four examples (G079.87+01.17, G074.56+00.84, G012.68-00.18, and G209.00-19.38) of $D$-$A$ diagrams toward the direction of masers. The distance of the last jump point agrees well with masers.

### v2

@@ A.2 The “noise” of CUSUM
$E\left[\left|S_{n}\right|\right]=\int_{-\infty}^{\infty}|x|\frac{1}{\sqrt{2\pi n\sigma^{2}}}e^{-\frac{(x-n\mu)^{2}}{2n\sigma^{2}}}dx,$ (A4)
where $\mu$ is the absolute difference between the mean of each segment and the overall mean $\bar{A}$, $\sigma$ is the dispersion of $A_{\lambda}$ of each segment, and $n$ is the equivalent number of steps. This expression does not have a simple elementary function form, but it can be split into two parts and solved using the error function ($\rm erf$), which is a related function of the Gaussian integral,
@@
The intricate statistical relationship between $S^{0}_{\rm diff}$ and $S_{\rm diff}$ allows for more effective utilization of physical quantities such as $\Delta A_{\lambda}$, $\sigma$, and $n$, aligning closely with outcomes from manual verification. Moreover, validations and implementations in this work demonstrate that the derived “noise” in CUSUM offers an efficient approach to assess the $CL$.
Figure A2: The $D$-$S$ plot in the presence of a slope in $D$-AA space. Six false change points are noted by red crosses (I, II, III, IV, V, VI are the serial numbers of the detection in iterations). The curve is best fit by using the least squares method (in green). In comparison, a predicted parabola curve (in blue) is drawn using our preset parameters. The colored curves are the $D$-$S$ of shuffled samples.
@@ A.3 The False Jump Features from Slope in D-A Map
We assume the extinction of diffuse gas still follows a Gaussian distribution on the $D$-AA map, yet it is superimposed on a linear function. The CUSUM statistic $S$ satisfies
@@
In the case, we set $S\left(0\right)=0$. The location of the symmetry axis, and/or the first false change point to be detected, is in $x_{\rm I}=\frac{\left(\bar{Y}-m\right)}{k}+\frac{1}{2n}$. Because $\frac{1}{2n}$ is smaller than the distance resolution in a given number density $n$, we ignore it and $x_{\rm I}\approx\frac{\left(\bar{Y}-m\right)}{k}$. CUSUM $S$ can thus be written as
@@
As illustrated in Figure A2, the $D$-$S$ plot draws a parabolic trace as expected. In a smaller $CL$ threshold (CC=1 in Mode I), six false change points were detected in order, all of which are located near the midpoints of their respective iteration intervals. In comparison, we give the best fit by using the least squares method (green curve in Figure A2) and a predicted parabola curve from the preset parameters (blue curve in Figure A2) by Equation (A10). A stricter criterion is needed to avoid such false detections. Next, we have to further analyze the properties of “noise” $S_{\rm diff}$.
Figure A3: Relations between $S_{\rm diff}$ and $A_{\lambda}$ dispersion $\sigma$, number density $n$, and slope $k$. (a) Set $n=0.1$ ${\rm pc}^{-1}$ and $k=0.003$ ${\rm mag}\cdot{\rm pc}^{-1}$, $S^{\rm avg}_{\rm diff}\propto\sigma$ when $\sigma$ becomes larger. (b) $S^{\rm avg}_{\rm diff}\propto\sqrt{n}$. (c) Set $\sigma$=1 mag, the relation between $S^{\rm avg}_{\rm diff}$ and $k$ shows a similar trend to panel (a). Panel (d) is a zoom-in from panel (c).
Similar to the situation where there are jump points, the $S_{\rm diff}$ with a slope of $k$ and dispersion $\sigma$ follows a random walk with variable steps (see Appendix A.1). Providing an analytical form for $S_{\rm diff}$ is equally complex. Generally, $S_{\rm diff}$ is proportional to $\sqrt{n}$, while $k$ and $\sigma$ also present a competitive relationship. For example, when $k$ is small, $\sigma$ becomes dominant. The relation between $S_{\rm diff}$, $n$, $k$, and $\sigma$ can also be derived using mock data. By examining the typical values from the $D$-$S$ diagram along the LoS, we can describe the relationship between $S_{\rm diff}$ and one of the quantities.
@@
However, we find that $\sigma$ domains $S_{\rm diff}$ in small $k$ values. Figure A3 (d) shows a zoomed-in view of the small $k$ region, where $S_{\rm diff}$ climbs slowly within the range of small $k$. In a simple test of the mock data, assuming the following preset parameters: a typical value of jumps intervals $x_{\rm I}=750$ pc, a gradient $k=0.002$ ${\rm mag}\cdot{\rm pc}^{-1}$, a stars density $n=0.1~{\rm pc}^{-1}$, and $A_{\lambda}$ dispersion $\sigma=1~{\rm mag}$, we compute the prominence of the preset slope. The calculated prominence is $S_{\rm diff}^{0}\sim 50$ and the averaged $S_{\rm diff}\sim 20$. This can effectively avoid false detections caused by slopes when setting $C\geq 2$, and the percentile to 0.95. Of course, for larger $\Delta A_{\lambda}$, we could raise the confidence threshold to ensure a more reliable result. On the other hand, if $k\gtrsim 0.003$ ${\rm mag}\cdot{\rm pc}^{-1}$, we can find these change points near midpoint of each iteration, due to the symmetric/parabolic nature of $S_{\rm diff}$. This may help us determine if a extinction structure is from a real jump or a large slope $k$ (see Figure B1 in Appendix B).
@@ Appendix C Examples for HMSFRs associated with masers
Figure C1: Four examples for G079.87+01.17, G074.56+00.84, G012.68-00.18, and G209.00-19.38. Stars are represented by gray dots, while black dots denote binned points every 20 pc. All four examples corresponding to the masers are located at the last jump point.
In Section 4.2, we found that a total of 46 masers align well with our results. Here, we provide four examples (G079.87+01.17, G074.56+00.84, G012.68-00.18, and G209.00-19.38) of $D$-AA diagrams (see Figure C1) toward the direction of masers. The distance of the last jump point agrees well with the masers.